A water pipe narrows from a radius of r1 = 4.99 cm to a radius of r2 = 2.75 cm. If the speed of the water in the wider part of the pipe is 1.31 m/s, what is the speed of the water in the narrower part? 



1
Expert's answer
2021-12-12T16:42:10-0500

"v_1S_1=v_2S_2,"

"v_1r_1^2=v_2r_2^2,"

"v_1=v_2\frac{r_2^2}{r_2^2}=4.31~\frac ms."